OpenStudy (anonymous):

(X-2)^2+(X-2)-12=0 solve using appropriate technique?????

OpenStudy (anonymous):

consider x-2 as y and the equation becomes y^2+y-12=0.... implying y=3,4.. using the standard method to sove a quadratic equation... and hence x-2=3 or x-2=4, impying x=5 or x=6.

OpenStudy (anonymous):

thank you